# Euler-type approximation for the invariant measure: An abstract framework

### v1

@@ Appendix B Additional technical proofs for some applications
In contrast with the $W_{1}$ case (see Theorem 5.6), we are not able to deduce the coupling property from the sewing lemma, see Eq. (3.1). Thus, we prove here directly the coupling property. To construct the coupling, we work with approximation schemes on a regular time grid and let then the time step go to zero to obtain the coupling property for the flow. Namely, for each $m\in{\mathbb{N}}$ we consider the grid
@@
In order to prove Theorem 5.7, we use Theorem 2.15 so we have to produce a probabilistic representation and to prove that this representation verifies (2.22) and (3.4). For the moment we fix $m$, and for $s_{i-1}\leq u<s_{i}$ we denote $\eta(u)=s_{i-1}$.
@@
We construct now a joint probabilistic representation (in the sense defined in Section 1). We consider a filtered probability space $(\Omega,\mathcal{F},\mathbb{P})$ that endows $dN(w,z,r)$ a Poisson point measure of intensity $dw\nu(dz)dr$. We denote by $(\mathcal{F}_{t})_{t\geq 0}$ the filtration generated by this random measure and $d\tilde{N}$ the compensated measure. Moreover, we consider random variables $X^{i}\in L^{2}(\Omega),i=1,2$ which are $\mathcal{F}_{s}$ measurable and such that $(X^{1},X^{2})$ has the law $\Pi_{s,s}(dv_{1},dv_{2}),$ which is the $W_{2}$ optimal coupling of $\mu^{1}$ and $\mu^{2}.$ In particular $X^{i}\sim\mu^{i}$. Then, we define the equations which give the coupling
@@
Notice first that $\Theta_{s,t}^{m}(\mu^{1})$ is the law of $X_{s,t}^{1}(X^{1})$ constructed recursively from (5.34) by
$X_{s,t}^{1}(X^{1})$ $=$ $X^{1}+\int_{s}^{t}b(x_{s,\eta(r)}^{1}(X^{1}))dr+\sum_{s_{j}\leq t}\int_{s_{j-1}}^{s_{j}}\int_{E}\int_{\mathbb{R}^{d}}c(v,z,X_{s,s_{j-1}}^{1}(X^{1}))d\widetilde{N}_{\rho_{r}}(v,z,r)$
@@
$\lambda_{h}(x)$ $=$ $\int_{0}^{1}\int_{E}\left(e^{i\left\langle h,c(\tau_{s,s_{j-1}}^{1}(w),z,x)\right\rangle}-1-i\langle h,c(\tau_{s,s_{j-1}}^{1}(w),z,x)\rangle\right)\nu(dz)dw$
$\lambda_{h}^{\prime}(x)$ $=$ $\int_{\mathbb{R}^{d}}\int_{E}\left(e^{i\langle h,c(v,z,x)\rangle}-1-i\langle h,c(v,z,x)\rangle\right)\nu(dz)\Theta_{s,s_{j-1}}^{m}(\mu^{1})(dv).$
@@
Both $x_{s,t}^{1}(X^{1})$ and $x_{s,t}^{2}(X^{2})$ are defined on the same probability space - so they represent a joint probabilistic representation. But we have to check the hypothesis concerning the coupling: see (2.21), (2.22 ),(3.4). By the very construction (2.21) is verified.
@@
From the assumptions (5.35) and (5.36), we have for every $\delta>0$
$\left\langle x,b(x)\right\rangle$ $\leq$ $(-\overline{b}+\delta)\left|x\right|^{2}+\delta^{-1}\left|b(0)\right|^{2}$ (B.15)
$\left|c(v,z,x)\right|^{2}$ $\leq$ $\overline{c}^{2}(z)(1+\left|v\right|^{2}+\left|x\right|^{2})\quad and\quad\left|b(x)\right|\leq K(1+\left|x\right|).$ (B.16)
Lemma B.5.
Under hypotheses (5.35) and (5.36), we have the following:
@@
${\mathbb{E}}\left(|x_{s,t}^{i}(X^{i})-X^{i}|^{2}\right)\leq C(1+{\mathbb{E}}\left(|X^{i}|^{2}\right))(t-s).$ (B.17)
@@ B If $\overline{b}>Q+\delta$ then for every $s<t\leq s+1$
${\mathbb{E}}\left(|x_{s,t}^{1}|^{2}\right)$ $\leq$ ${\mathbb{E}}\left(|X^{1}|^{2}\right)(1-2(\overline{b}-Q-\delta)(t-s))$
$+(2\delta^{-1}\left|b(0)\right|^{2}+Q)(t-s)+C(1+{\mathbb{E}}\left(|X^{1}|^{2}\right))(t-s)^{3/2}.$
Note that the result in (B.5) differs in rate from the one in Lemma B.1.
@@
We start from the equation satisfied by $x_{s,t}^{1}$ in (B.14). Using the inequality $|x+y|^{2}\leq 2(|x|^{2}+|y|^{2})$, Cauchy-Schwarz inequality, the $L^{2}$-isometry for martingale driven stochastic integrals, (5.38), (B.15) and (B.16), one gets
${\mathbb{E}}\left(|x_{s,t}^{1}|^{2}\right)\leq C\left(1+{\mathbb{E}}\left|X^{1}\right|^{2}+\int_{s}^{t}{\mathbb{E}}\left(|x_{s,\eta(r)}^{1}|^{2}\right)dr\right).$
Here, we have used that $\int_{0}^{1}|\tau_{s,\eta(r)}^{1}(w)|^{2}dw={\mathbb{E}}(|x_{s,\eta(r)}^{1}|^{2})$ and the constant $C$ depends on $Q$. This identity will be used frequently without any further mention of it. Next, due to a Gronwall type inequality we obtain
${\mathbb{E}}(|x_{s,t}^{1}|^{2})\leq C(1+{\mathbb{E}}\left|X^{1}\right|^{2}).$ (B.19)
Proof of A: Let us prove (B.17). As before, note that using the inequality $|x+y|^{2}\leq 2(|x|^{2}+|y|^{2})$, Cauchy-Schwarz inequality, the $L^{2}$-isometry for martingale driven stochastic integrals, (B.16) and (B.19) one obtains
${\mathbb{E}}\left(|x_{s,t}^{1}(X^{1})-X^{1}|^{2}\right)$
$\leq$ $2(t-s){\mathbb{E}}\int_{s}^{t}|b(x_{s,\eta(r)}^{1})|^{2}dr+2{\mathbb{E}}\int_{s}^{t}\int_{E}\int_{0}^{1}|c(\tau_{s,\eta(r)}^{1}(w),z,x_{s,\eta(r)}^{1})|^{2}dwd\nu(z)dr$
$\leq$ $2K(t-s){\mathbb{E}}\int_{s}^{t}(1+|x_{s,\eta(r)}^{1}|^{2})dr+2Q\int_{s}^{t}(1+2{\mathbb{E}}(|x_{s,\eta(r)}^{1}|^{2}))dr$
$\leq$ $C^{\prime}(t-s)(1+{\mathbb{E}}\left(|X^{1}|^{2}\right)).$
@@
Proof of B: One employs (B.17) and (B.15) to obtain
${\mathbb{E}}\left(\langle x_{s,r}^{1},b(x_{s,\eta(r)}^{1})\rangle\right)$ $\leq$ $\left({\mathbb{E}}(|x_{s,r}^{1}-x_{s,\eta(r)}^{1}|^{2}){\mathbb{E}}(|b(x_{s,\eta(r)}^{1})|^{2})\right)^{1/2}+{\mathbb{E}}\left(\langle x_{s,\eta(r)}^{1},b(x_{s,\eta(r)}^{1})\rangle\right)$
$\leq$ $C(r-\eta(r))^{1/2}(1+{\mathbb{E}}(|X^{1}|^{2}))+(-\overline{b}+\delta){\mathbb{E}}(|x_{s,\eta(r)}^{1}|^{2})+\delta^{-1}|b(0)|^{2}$
Since $\int_{0}^{1}\left|\tau_{s,\eta(r)}^{1}(w)\right|^{2}dw={\mathbb{E}}\left|x_{s,\eta(r)}^{1}\right|^{2}$ and we have assumed ( 5.38) we also have
${\mathbb{E}}\left(\int_{s}^{t}\int_{E}\int_{0}^{1}\overline{c}^{2}(z)(1+|\tau_{s,\eta(r)}^{1}(w)|^{2}+|x_{s,\eta(r)}^{1}|^{2})dwd\nu(z)dr\right)\leq Q\int_{s}^{t}(1+2{\mathbb{E}}(|x_{s,\eta(r)}^{1}|^{2}))dr$
@@
${\mathbb{E}}\left(|x_{s,t}^{1}|^{2}\right)=$ ${\mathbb{E}}(|X^{1}|^{2})+2{\mathbb{E}}\int_{s}^{t}\left\langle x_{s,r}^{1},b(x_{s,\eta(r)}^{1})\right\rangle dr$
$+{\mathbb{E}}\int_{s}^{t}\int_{E}\int_{0}^{1}\left|c(\tau_{s,\eta(r)}^{1}(w),z,x_{s,\eta(r)}^{1})\right|^{2}dwd\nu(z)dr.$
$\leq$ ${\mathbb{E}}\left(|X^{1}|^{2}\right)+C(t-s)^{3/2}(1+{\mathbb{E}}(|X^{1}|^{2}))-2(\overline{b}-Q-\delta){\mathbb{E}}\int_{s}^{t}|x_{s,\eta(r)}^{1}|^{2}dr$
@@
We have $|{\mathbb{E}}(|x_{s,\eta(r)}^{1}|^{2})-{\mathbb{E}}(|X^{1}|^{2})|\leq\sqrt{{\mathbb{E}}((|x_{s,\eta(r)}^{1}|+|X^{1}|)^{2})}\sqrt{{\mathbb{E}}(|x_{s,\eta(r)}^{1}-X^{1}|^{2})}$, and we get by using A and (B.19)
$\int_{s}^{t}\left|{\mathbb{E}}(|x_{s,\eta(r)}^{1}|^{2})-{\mathbb{E}}(|X^{1}|^{2})\right|dr\leq C(1+{\mathbb{E}}\left|X^{1}\right|^{2})(t-s)^{3/2}.$
@@
${\mathbb{E}}(|x_{s,t}^{1}|^{2})$ $\leq$ $(1-2(\overline{b}-Q-\delta)(t-s)){\mathbb{E}}\left|X^{1}\right|^{2}$
$+C(t-s)^{3/2}(1+{\mathbb{E}}(|X^{1}|^{2}))+(2\delta^{-1}\left|b(0)\right|^{2}+Q)(t-s).\qed$
We conclude from Lemma B.5 that
@@
We remark that the difference in the value of $\varepsilon^{\prime}$ between (B.5) and the result in Lemma B.1 is due to the fact that the above proof uses the modulus of continuity of the underlying process. This difference will also appear in the rest of the proof.
@@
Lemma B.6.
A We assume that assumptions (5.35) and (5.36) hold with $Q<\overline{b}.$ Then, for every $b_{*}<2(\overline{b}-Q)$ there exists a constant $C$ such that, if $0\leq t-s\leq 1,$ then
${\mathbb{E}}\left|x_{s,t}^{1}(X^{1})-x_{s,t}^{2}(X^{2})\right|^{2}\leq{\mathbb{E}}\left|X^{1}-X^{2}\right|^{2}(1-b_{*}(t-s))+C(1+{\mathbb{E}}\left|X^{1}\right|^{2}+{\mathbb{E}}\left|X^{2}\right|^{2})(t-s)^{3/2}.$ (B.20)
@@ B Moreover $\theta$ is $(2,b_{*},1)$-self-coupled for some $h_{0}>0$ and more precisely
$W^{2}_{2}(\theta_{s,t}(\mu^{1}),\theta_{s,t}(\mu^{2}))\leq(1-b_{*}(t-s))W^{2}_{2}(\mu^{1},\mu^{2}){+C(1+\|\mu_{1}\|^{2}_{2}+\|\mu_{2}\|^{2})(t-s)^{2}}.$ (B.21)
Remark B.7.
When dealing with $W_{2}(\theta_{s,t}(\mu^{1}),\theta_{s,t}(\mu^{2}))$ no error terms of the type ${\mathbb{E}}\left|b(x_{s,r}^{2}(X^{2}))-b(X^{2})\right|^{2}+{\mathbb{E}}\left|b(x_{s,\eta(r)}^{1})-b(x_{s,r}^{1})\right|^{2}$ appear in the proof. This is the reason why the term $C(1+{\mathbb{E}}\left|X^{1}\right|^{2}+{\mathbb{E}}\left|X^{2}\right|^{2})(t-s)^{3/2}$ does not appear in the final result in comparison with (B.20).
@@
As in the previous proof we simplify the notation using $(x_{s,t}^{1},x_{s,t}^{2})=(x_{s,t}^{1}(X^{1}),x_{s,t}^{2}(X^{2}))$. Proof of A. The process $y_{s,t}=x_{s,t}^{1}-x_{s,t}^{2}$ satisfies
@@
$\Delta c(r,w,z)$ $=$ $c(\tau_{s,\eta(r)}^{1}(w),z,x_{s,\eta(r)}^{1})-c(\tau_{s,\eta(r)}^{2}(w),z,X^{2})$
From (5.35) and using Lemma B.5, we get for every $\delta>0$
${\mathbb{E}}\left\langle y_{s,r},\Delta b(r)\right\rangle$ $\leq$ $-\overline{b}{\mathbb{E}}(|y_{s,r}|^{2})+{\mathbb{E}}\left\langle y_{s,r},b(x_{s,\eta(r)}^{1})-b(x_{s,r}^{1})\right\rangle+{\mathbb{E}}\left\langle y_{s,r},b(x_{s,r}^{2})-b(X^{2})\right\rangle$
$\leq$ $-\overline{b}{\mathbb{E}}(|y_{s,r}|^{2})+2\delta{\mathbb{E}}(|y_{s,r}|^{2})$
$+{\delta^{-1}}\left({\mathbb{E}}\left(|b(x_{s,r}^{2})-b(X^{2})|^{2}\right)+{\mathbb{E}}\left(|b(x_{s,\eta(r)}^{1})-b(x_{s,r}^{1})|^{2}\right)\right)$
$\leq$ $(2\delta-\overline{b}){\mathbb{E}}\left|y_{s,r}\right|^{2}+{\delta^{-1}}C_{b}\left({\mathbb{E}}\left|x_{s,r}^{2}-X^{2}\right|^{2}+{\mathbb{E}}\left|x_{s,\eta(r)}^{1}-x_{s,r}^{1}\right|^{2}\right)$
$\leq$ $(2\delta-\overline{b}){\mathbb{E}}\left|y_{s,r}\right|^{2}+{\delta^{-1}}C_{b}C(1+{\mathbb{E}}\left|X^{1}\right|^{2}+{\mathbb{E}}\left|X^{2}\right|^{2})(r-s).$
Using the inequality $\left|\left|a\right|^{2}-\left|b\right|^{2}\right|\leq\left|a-b\right|(\left|a\right|+\left|b\right|)$, Cauchy-Schwarz inequality and (B.17), we have
$\left|{\mathbb{E}}\left|y_{s,r}\right|^{2}-{\mathbb{E}}\left|X^{1}-X^{2}\right|^{2}\right|\leq C(1+{\mathbb{E}}\left|X^{1}\right|^{2}+{\mathbb{E}}\left|X^{2}\right|^{2})(r-s)^{1/2}.$ (B.22)
@@
${\mathbb{E}}\left\langle y_{s,r},\Delta b(r)\right\rangle\leq(2\delta-\overline{b}){\mathbb{E}}\left|X^{1}-X^{2}\right|^{2}+C(1+{\mathbb{E}}\left|X^{1}\right|^{2}+{\mathbb{E}}\left|X^{2}\right|^{2})(r-s)^{\frac{1}{2}}$
We now focus on the jump term and have by (5.36)

### v2

@@ C.2. A Boltzmann equation in the martingale regime ($W_{2}$)
This appendix section is devoted to the proof of this theorem, which follows the same steps as the one with $W_{1}$ estimates. However, in contrast with the $W_{1}$ case (see Theorem 5.5), we are not able to deduce the coupling property from the sewing lemma, see Eq. (3.1). Thus, we prove here directly the coupling property. To construct the coupling, we work with approximation schemes on a regular time grid and let then the time step go to zero to obtain the coupling property for the flow. Namely, for each $m\in{\mathbb{N}}$ we consider the grid
@@
In order to prove Theorem C.4, we use Theorem 2.19 so we have to produce a probabilistic representation and to prove that this representation verifies (2.27) and (3.4). For the moment we fix $m$, and for $s_{i-1}\leq u<s_{i}$ we denote $\eta(u)=s_{i-1}$.
@@
We construct now a joint probabilistic representation (in the sense defined in Section 1). We consider a filtered probability space $(\Omega,\mathcal{F},\mathbb{P})$ that endows $dN(w,z,r)$ a Poisson point measure of intensity $dw\Lambda(dz)dr$. We denote by $(\mathcal{F}_{t})_{t\geq 0}$ the filtration generated by this random measure and $d\tilde{N}$ the compensated measure. Moreover, we consider random variables $X^{i}\in L^{2}(\Omega),i=1,2$ which are $\mathcal{F}_{s}$ measurable and such that $(X^{1},X^{2})$ has the law $\Pi_{s,s}(dv_{1},dv_{2}),$ which is the $W_{2}$ optimal coupling of $\mu^{1}$ and $\mu^{2}.$ In particular $X^{i}\sim\mu^{i}$. Then, we define the equations which give the coupling
@@
Notice first that $\Theta_{s,t}^{m}(\mu^{1})$ is the law of $X_{s,t}^{1}(X^{1})$ constructed recursively from (C.11) by
$X_{s,t}^{1}(X^{1})=$ $X^{1}+\int_{s}^{t}b(x_{s,\eta(r)}^{1}(X^{1}))dr$
$+\sum_{s_{j}\leq t}\int_{s_{j-1}}^{s_{j}}\int_{E}\int_{\mathbb{R}^{d}}c(v,z,X_{s,s_{j-1}}^{1}(X^{1}))d\widetilde{N}_{\rho_{r}}(v,z,r)$
@@
$\lambda_{h}(x)$ $=\int_{0}^{1}\int_{E}\left(e^{i\left\langle h,c(\tau_{s,s_{j-1}}^{1}(w),z,x)\right\rangle}-1-i\langle h,c(\tau_{s,s_{j-1}}^{1}(w),z,x)\rangle\right)\Lambda(dz)dw$
$\lambda_{h}^{\prime}(x)$ $=\int_{\mathbb{R}^{d}}\int_{E}\left(e^{i\langle h,c(v,z,x)\rangle}-1-i\langle h,c(v,z,x)\rangle\right)\Lambda(dz)\Theta_{s,s_{j-1}}^{m}(\mu^{1})(dv).$
@@
Both $x_{s,t}^{1}(X^{1})$ and $x_{s,t}^{2}(X^{2})$ are defined on the same probability space - so they represent a joint probabilistic representation. But we have to check the hypothesis concerning the coupling: see (2.26), (2.27 ),(3.4). By the very construction (2.26) is verified.
@@
From the assumptions (C.12) and (C.13), we have for every $\delta>0$
$\left\langle x,b(x)\right\rangle$ $\leq(-\overline{b}+\delta)\left|x\right|^{2}+\delta^{-1}\left|b(0)\right|^{2}$ (C.18)
$\left|c(v,z,x)\right|^{2}$ $\leq\overline{c}^{2}(z)(1+\left|v\right|^{2}+\left|x\right|^{2})\quad and\quad\left|b(x)\right|\leq K(1+\left|x\right|).$ (C.19)
Lemma C.6.
Under hypotheses (C.12) and (C.13), we have the following:
@@
${\mathbb{E}}[|x_{s,t}^{i}(X^{i})-X^{i}|^{2}]\leq C(1+{\mathbb{E}}[|X^{i}|^{2}])(t-s).$ (C.20)
@@ B If $\overline{b}>Q+\delta$ then for every $s<t\leq s+1$
${\mathbb{E}}[|x_{s,t}^{1}|^{2}]\leq$ ${\mathbb{E}}[|X^{1}|^{2}](1-2(\overline{b}-Q-\delta)(t-s))$ (C.21)
$+(2\delta^{-1}\left|b(0)\right|^{2}+Q)(t-s)+C(1+{\mathbb{E}}[|X^{1}|^{2}])(t-s)^{3/2}.$
Note that the result in (C.21) differs in rate from the one in Lemma C.1.
@@
We start from the equation satisfied by $x_{s,t}^{1}$ in (C.17). Using the inequality $|x+y|^{2}\leq 2(|x|^{2}+|y|^{2})$, Cauchy-Schwarz inequality, the $L^{2}$-isometry for martingale driven stochastic integrals, (C.15), (C.18) and (C.19), one gets
${\mathbb{E}}[|x_{s,t}^{1}|^{2}]\leq C\left(1+{\mathbb{E}}[|X^{1}|^{2}]+\int_{s}^{t}{\mathbb{E}}\left[|x_{s,\eta(r)}^{1}|^{2}\right]dr\right).$
Here, we have used that $\int_{0}^{1}|\tau_{s,\eta(r)}^{1}(w)|^{2}dw={\mathbb{E}}[|x_{s,\eta(r)}^{1}|^{2}]$ and the constant $C$ depends on $Q$. This identity will be used frequently without any further mention of it. Next, due to a Gronwall type inequality we obtain
${\mathbb{E}}[|x_{s,t}^{1}|^{2}]\leq C(1+{\mathbb{E}}[|X^{1}|^{2}]).$ (C.22)
Proof of A: Let us prove (C.20). As before, note that using the inequality $|x+y|^{2}\leq 2(|x|^{2}+|y|^{2})$, Cauchy-Schwarz inequality, the $L^{2}$-isometry for martingale driven stochastic integrals, (C.19) and (C.22) one obtains
${\mathbb{E}}\left[|x_{s,t}^{1}(X^{1})-X^{1}|^{2}\right]$
$\leq 2(t-s){\mathbb{E}}\int_{s}^{t}|b(x_{s,\eta(r)}^{1})|^{2}dr+2{\mathbb{E}}\int_{s}^{t}\int_{E}\int_{0}^{1}|c(\tau_{s,\eta(r)}^{1}(w),z,x_{s,\eta(r)}^{1})|^{2}dw\Lambda(dz)dr$
$\leq 2K(t-s){\mathbb{E}}\int_{s}^{t}(1+|x_{s,\eta(r)}^{1}|^{2})dr+2Q\int_{s}^{t}(1+2{\mathbb{E}}[|x_{s,\eta(r)}^{1}|^{2}])dr$
$\leq C^{\prime}(t-s)(1+{\mathbb{E}}[|X^{1}|^{2}]).$
@@
Proof of B: One employs (C.20) and (C.18) to obtain
${\mathbb{E}}\left[\langle x_{s,r}^{1},b(x_{s,\eta(r)}^{1})\rangle\right]\leq$ $\left({\mathbb{E}}[|x_{s,r}^{1}-x_{s,\eta(r)}^{1}|^{2}]{\mathbb{E}}[|b(x_{s,\eta(r)}^{1})|^{2}]\right)^{1/2}$
$+{\mathbb{E}}\left[\langle x_{s,\eta(r)}^{1},b(x_{s,\eta(r)}^{1})\rangle\right]$
$\leq$ $C(r-\eta(r))^{1/2}(1+{\mathbb{E}}[|X^{1}|^{2}])+(-\overline{b}+\delta){\mathbb{E}}[|x_{s,\eta(r)}^{1}|^{2}]$
$+\delta^{-1}|b(0)|^{2}.$
Since $\int_{0}^{1}\left|\tau_{s,\eta(r)}^{1}(w)\right|^{2}dw={\mathbb{E}}[(x_{s,\eta(r)}^{1})^{2}]$ and we have assumed ( C.15) we also have
${\mathbb{E}}\left[\int_{s}^{t}\!\!\int_{E}\int_{0}^{1}\!\!\overline{c}^{2}(z)(1+|\tau_{s,\eta(r)}^{1}(w)|^{2}+|x_{s,\eta(r)}^{1}|^{2})dw\Lambda(dz)dr\!\right]\!\leq Q\!\!\int_{s}^{t}\!\!(1+2{\mathbb{E}}[|x_{s,\eta(r)}^{1}|^{2}])dr.$
@@
${\mathbb{E}}\left[|x_{s,t}^{1}|^{2}\right]=$ ${\mathbb{E}}[|X^{1}|^{2}]+2{\mathbb{E}}\int_{s}^{t}\left\langle x_{s,r}^{1},b(x_{s,\eta(r)}^{1})\right\rangle dr$
$+{\mathbb{E}}\int_{s}^{t}\int_{E}\int_{0}^{1}\left|c(\tau_{s,\eta(r)}^{1}(w),z,x_{s,\eta(r)}^{1})\right|^{2}dw\Lambda(dz)dr.$
$\leq$ ${\mathbb{E}}[|X^{1}|^{2}]+C(t-s)^{3/2}(1+{\mathbb{E}}[|X^{1}|^{2}])-2(\overline{b}-Q-\delta){\mathbb{E}}\int_{s}^{t}|x_{s,\eta(r)}^{1}|^{2}dr$
@@
We have $|{\mathbb{E}}[|x_{s,\eta(r)}^{1}|^{2}]-{\mathbb{E}}[|X^{1}|^{2}]|\leq\sqrt{{\mathbb{E}}[(|x_{s,\eta(r)}^{1}|+|X^{1}|)^{2}]}\sqrt{{\mathbb{E}}[|x_{s,\eta(r)}^{1}-X^{1}|^{2}]}$, and we get by using A and (C.22)
$\int_{s}^{t}\left|{\mathbb{E}}[|x_{s,\eta(r)}^{1}|^{2}]-{\mathbb{E}}[|X^{1}|^{2}]\right|dr\leq C(1+{\mathbb{E}}[|X^{1}|^{2}])(t-s)^{3/2}.$
@@
${\mathbb{E}}[|x_{s,t}^{1}|^{2}]\leq$ $(1-2(\overline{b}-Q-\delta)(t-s)){\mathbb{E}}[|X^{1}|^{2}]$
$+C(t-s)^{3/2}(1+{\mathbb{E}}[|X^{1}|^{2}])+(2\delta^{-1}\left|b(0)\right|^{2}+Q)(t-s).\qed$
We conclude from Lemma C.6 that
@@
We remark that the difference in the value of $\varepsilon^{\prime}$ between (C.21) and the result in Lemma C.1 is due to the fact that the above proof uses the modulus of continuity of the underlying process. This difference will also appear in the rest of the proof.
@@
Lemma C.7.
A We assume that assumptions (C.12) and (C.13) hold with $Q<\overline{b}.$ Then, for every $b_{*}<2(\overline{b}-Q)$ there exists a constant $C$ such that, if $0\leq t-s\leq 1,$ then
${\mathbb{E}}[|x_{s,t}^{1}(X^{1})-x_{s,t}^{2}(X^{2})|^{2}]\leq$ ${\mathbb{E}}[|X^{1}-X^{2}|^{2}](1-b_{*}(t-s))+C(1+{\mathbb{E}}[|X^{1}|^{2}]$
$+{\mathbb{E}}[|X^{2}|^{2}])(t-s)^{3/2}.$ (C.23)
@@ B Moreover $\theta$ is $(2,b_{*},1)$-self-coupled for some $h_{0}>0$ and more precisely
$W^{2}_{2}(\theta_{s,t}(\mu^{1}),\theta_{s,t}(\mu^{2}))\leq(1-b_{*}(t-s))W^{2}_{2}(\mu^{1},\mu^{2}){+C(1+\|\mu_{1}\|^{2}_{2}+\|\mu_{2}\|^{2})(t-s)^{2}}.$ (C.24)
Remark C.8.
When dealing with $W_{2}(\theta_{s,t}(\mu^{1}),\theta_{s,t}(\mu^{2}))$ no error terms of the type ${\mathbb{E}[}|b(x_{s,r}^{2}(X^{2}))-b(X^{2})|^{2}]+{\mathbb{E}}[|b(x_{s,\eta(r)}^{1})-b(x_{s,r}^{1})|^{2}]$ appear in the proof. This is the reason why the term $C(1+{\mathbb{E}}[|X^{1}|^{2}]+{\mathbb{E}}[|X^{2}|^{2}])(t-s)^{3/2}$ does not appear in the final result in comparison with (C.23).
@@
As in the previous proof we simplify the notation using $(x_{s,t}^{1},x_{s,t}^{2})=(x_{s,t}^{1}(X^{1}),x_{s,t}^{2}(X^{2}))$.
Proof of A. The process $y_{s,t}=x_{s,t}^{1}-x_{s,t}^{2}$ satisfies
@@
$\Delta c(r,w,z)$ $=c(\tau_{s,\eta(r)}^{1}(w),z,x_{s,\eta(r)}^{1})-c(\tau_{s,\eta(r)}^{2}(w),z,X^{2}).$
From (C.12) and using Lemma C.6, we get for every $\delta>0$
${\mathbb{E}}[\left\langle y_{s,r},\Delta b(r)\right\rangle]\leq$ $-\overline{b}{\mathbb{E}}[|y_{s,r}|^{2}]$
$+{\mathbb{E}}\left\langle y_{s,r},b(x_{s,\eta(r)}^{1})-b(x_{s,r}^{1})\right\rangle+{\mathbb{E}}\left\langle y_{s,r},b(x_{s,r}^{2})-b(X^{2})\right\rangle$
$\leq$ $-\overline{b}{\mathbb{E}}[|y_{s,r}|^{2}]+2\delta{\mathbb{E}}[|y_{s,r}|^{2}]$
$+{\delta^{-1}}\left({\mathbb{E}}\left[|b(x_{s,r}^{2})-b(X^{2})|^{2}\right]+{\mathbb{E}}\left[|b(x_{s,\eta(r)}^{1})-b(x_{s,r}^{1})|^{2}\right]\right)$
$\leq$ $(2\delta-\overline{b}){\mathbb{E}}[|y_{s,r}|^{2}]+{\delta^{-1}}C_{b}\left({\mathbb{E}}\left[|x_{s,r}^{2}-X^{2}|^{2}\right]+{\mathbb{E}}[|x_{s,\eta(r)}^{1}-x_{s,r}^{1}|^{2}]\right)$
$\leq$ $(2\delta-\overline{b}){\mathbb{E}}[|y_{s,r}|^{2}]+{\delta^{-1}}C_{b}C(1+{\mathbb{E}}[|X^{1}|^{2}]+{\mathbb{E}}[|X^{2}|^{2}])(r-s).$
Using the inequality $||a|^{2}-|b|^{2}|\leq|a-b|(|a|+|b|)$, Cauchy-Schwarz inequality and (C.20), we have
$|{\mathbb{E}}[|y_{s,r}|^{2}]-{\mathbb{E}}[|X^{1}-X^{2}|^{2}]|\leq C(1+{\mathbb{E}}[|X^{1}|^{2}]+{\mathbb{E}}[|X^{2}|^{2}])(r-s)^{1/2}.$ (C.25)
@@
${\mathbb{E}}\left\langle y_{s,r},\Delta b(r)\right\rangle\leq(2\delta-\overline{b}){\mathbb{E}}[|X^{1}-X^{2}|^{2}]+C(1+{\mathbb{E}}[|X^{1}|^{2}]+{\mathbb{E}}[|X^{2}|^{2}])(r-s)^{\frac{1}{2}}.$
We now focus on the jump term and have by (C.13)